OpenStudy (anonymous):

how do you simplify (x²-1)

OpenStudy (anonymous):

(x+1)(x-1)

OpenStudy (anonymous):

This is a differnce of squares: a^2-b^2=(a-b)(a+b)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You must be careful. If you are given a question like, $$\frac{x^2-1}{x+1}=\frac{(x-1)(x+1)}{x+1}=x - 1$$ For in this case you must stipulate that $$x+1\ne 0$$ so that you can actually do this simplification.